OpenStudy (anonymous):

The cost of producing pots is the initial investment, $175, plus $3.50 in materials per pot. Therefore, the function C(p) = 175 + 3.5p defines the cost, C, of producing a number of pots, p. To the nearest cent, how much will it cost to produce 125 pots? (Enter only the number without the dollar sign, such as 25.03 for twenty-five dollars and 3 cents.)

OpenStudy (imstuck):

Fill 125 in for p in your function and solve it. Can you do that?

OpenStudy (jgirl128):

c(p) = 175 + 3.5p first, plug in p. c(p) = 175 + 3.5(125) now, simplify and solve.

OpenStudy (anonymous):

C=$175+$3.5(125) C=$175+$437.50 C=$612.50 Therefore the cost to produce 125 pots is $612.50